# bibliography.bib

### v3

@@ 8.3 Proof of lemmas related to the Gibbs sampler of a Bayesian linear regression model
It follows from the de-initializing property of the Markov chain [roberts:rosenthal:2001, Example 3] that
@@
$||\mathcal{L}(\sigma^{2}_{n})-\mathcal{L}(\sigma^{{}^{\prime}2}_{n})||_{TV}$
$\leq E\left[||\mathcal{L}(W_{n}/G_{n}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1})-\mathcal{L}(W^{\prime}_{n}/G_{n}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1})||_{TV}\right]$ By Proposition 2.2 of [sixta]
@@
$=E\left[||\mathcal{L}\left(\frac{1}{G_{n}}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1}\right)-\mathcal{L}\left(\left(1+\frac{\Delta}{W_{n}}\right)\frac{1}{G_{n}}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1}\right)||_{TV}\right]$ By Proposition 2.1 of [sixta]
$\leq E\left[E\left[\sup_{x>0}\left\{1-\frac{\pi_{1/G_{n}}(x)}{\pi_{(1+\Delta/W_{n})/G_{n}}(x)}\right\}\bigm{|}Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1}\right]\right]$ By Lemma 6.16 of of [markovmixing]
@@
$||\mathcal{L}(\sigma^{2}_{n})-\mathcal{L}(\sigma^{{}^{\prime}2}_{n})||_{TV}$ $\leq E\left[\sup_{x>0}\left\{1-\frac{x^{-\alpha-1}e^{-1/x}}{(1+\Delta/W_{n})^{\alpha}x^{-\alpha-1}e^{-(1+\Delta/W_{n})/x}}\right\}\right]$
@@
Define $\alpha^{\prime}=(k+\upsilon)/2$ and $\beta^{\prime}=\upsilon_{0}c^{2}_{0}/2$. Note that
@@
$=\frac{\frac{1}{(\sigma^{2})^{(k+\upsilon_{0})/2+1}}\exp\left(-\frac{1}{2\sigma^{2}}(Y-X\beta)^{T}(Y-X\beta)-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})-\frac{\upsilon_{0}c_{0}^{2}}{2\sigma^{2}}\right)}{\frac{\beta^{\prime\alpha^{\prime}}}{\Gamma(\alpha^{\prime})}\frac{1}{(\sigma^{2})^{(k+\upsilon)/2+1}}e^{-\upsilon_{0}c^{2}_{0}/2\sigma^{2}}(2\pi)^{-q/2}\det(\Sigma_{\beta})^{-1/2}e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma^{-1}_{\beta}(\beta-\beta_{0})}}$
$=\frac{\exp\left(-\frac{1}{2\sigma^{2}}(Y-X\beta)^{T}(Y-X\beta)\right)}{\frac{\beta^{\prime\alpha^{\prime}}}{\Gamma(\alpha^{\prime})}(2\pi)^{-q/2}\det(\Sigma_{\beta})^{-1/2}}$
$\leq\frac{\Gamma(\alpha^{\prime})}{\beta^{\prime\alpha^{\prime}}}(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}$
By Lemma 3.2, if $K=\frac{(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{\int_{B}g(\beta,\sigma^{2})d(\beta\times\sigma^{2})}\frac{\Gamma(\alpha^{\prime})}{\beta^{\prime\alpha^{\prime}}}$, then the result follows. ∎
Remark.
For Real Data Example 6.1, we want to prove that $E[||(\beta_{n},\sigma^{2}_{n})-(\beta^{\prime}_{n},\sigma^{{}^{\prime}2}_{n})||_{1}]<\infty$ for $n\in\mathbb{N}$.
@@
Thus, we have established that $E[||(\sigma^{2}_{n},\beta_{n})||_{1}]<\infty$ and $E[||(\sigma^{{}^{\prime}2}_{n},\beta^{\prime}_{n})||_{1}]<\infty$. By the triangle inequality, $E[||(\sigma^{2}_{n},\beta_{n})-(\sigma^{{}^{\prime}2}_{n},\beta^{\prime}_{n})||_{1}]<\infty$.

### v4

@@ 8.6 Proof of lemmas related to the Gibbs sampler of a Bayesian linear regression model
It follows from the de-initializing property of the Markov chain [44, Example 3] that
@@
[…]
$\leq E\left[||\mathcal{L}(W_{n}/G_{n}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1})-\mathcal{L}(W^{\prime}_{n}/G_{n}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1})||_{TV}\right]$
By Proposition 2.2 of [51]
@@
$=E\left[||\mathcal{L}\left(\frac{1}{G_{n}}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1}\right)-\mathcal{L}\left(\left(1+\frac{\Delta}{W_{n}}\right)\frac{1}{G_{n}}\mid Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1}\right)||_{TV}\right]$
By Proposition 2.1 of [51]
$\leq E\left[E\left[\sup_{x>0}\left\{1-\frac{\pi_{1/G_{n}}(x)}{\pi_{(1+\Delta/W_{n})/G_{n}}(x)}\right\}\bigm|Z_{n},\sigma_{n-1}^{2},\sigma^{{}^{\prime}2}_{n-1}\right]\right]$
By Lemma 6.16 of of [33]
@@
[…]
@@
Define $\alpha^{\prime}=(k+\upsilon)/2-2$ and $\beta^{\prime}=\upsilon_{0}c^{2}_{0}/2$. Note that
@@
$=\frac{\frac{1}{(\sigma^{2})^{(k+\upsilon_{0})/2+1}}\exp\left(-\frac{1}{2\sigma^{2}}(Y-X\beta)^{T}(Y-X\beta)-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})-\frac{\upsilon_{0}c_{0}^{2}}{2\sigma^{2}}\right)}{\frac{\beta^{\prime\alpha^{\prime}}}{\Gamma(\alpha^{\prime})}\frac{1}{(\sigma^{2})^{(k+\upsilon)/2-2+1}}e^{-\upsilon_{0}c^{2}_{0}/2\sigma^{2}}(2\pi)^{-q/2}\det(\Sigma_{\beta})^{-1/2}e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma^{-1}_{\beta}(\beta-\beta_{0})}}$
$=\frac{(\sigma^{2})^{-2}\exp\left(-\frac{1}{2\sigma^{2}}(Y-X\beta)^{T}(Y-X\beta)\right)}{\frac{\beta^{\prime\alpha^{\prime}}}{\Gamma(\alpha^{\prime})}(2\pi)^{-q/2}\det(\Sigma_{\beta})^{-1/2}}$
The numerator of the last equality is proportional to an inverse gamma density function with $\alpha^{*}=1$ and $\beta^{*}=\frac{1}{2}(Y-X\beta)^{T}(Y-X\beta)$. Thus, the mode occurs when $\sigma^{2*}=\frac{\beta^{*}}{\alpha^{*}+1}=\frac{(Y-X\beta)^{T}(Y-X\beta)}{4}$ and so,
$\leq\frac{16}{((Y-X\beta)^{T}(Y-X\beta))^{2}}\frac{\exp\left(-\frac{4}{2(Y-X\beta)^{T}(Y-X\beta)}(Y-X\beta)^{T}(Y-X\beta)\right)}{\frac{\beta^{\prime\alpha^{\prime}}}{\Gamma(\alpha^{\prime})}(2\pi)^{-q/2}\det(\Sigma_{\beta})^{-1/2}}$
$=\frac{16}{((Y-X\beta)^{T}(Y-X\beta))^{2}}\frac{\Gamma(\alpha^{\prime})(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{\beta^{\prime\alpha^{\prime}}e^{2}}$
$\leq\frac{16}{((Y-X\hat{\beta})^{T}(Y-X\hat{\beta}))^{2}}\frac{\Gamma(\alpha^{\prime})(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{\beta^{\prime\alpha^{\prime}}e^{2}}$
Where $\hat{\beta}=(X^{T}X)^{-1}X^{T}Y$ is the maximum likelihood estimator. Next we find a lower bound on $\int_{B}g(\beta,\sigma^{2})\mathrm{d}(\beta,\sigma^{2})$
$\int_{\mathbb{R}^{q}\times\mathbb{R}_{+}}g(\beta,\sigma^{2})\mathrm{d}(\beta,\sigma^{2})$
$=\int_{\mathbb{R}^{q}\times\mathbb{R}_{+}}\frac{1}{(\sigma^{2})^{(k+\upsilon_{0})/2+1}}$
$\hskip 56.9055pt\exp\left(-\frac{1}{2\sigma^{2}}(Y-X\beta)^{T}(Y-X\beta)-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})-\frac{\upsilon_{0}c_{0}^{2}}{2\sigma^{2}}\right)\mathrm{d}(\beta,\sigma^{2})$
$=\int_{\mathbb{R}^{q}}\left(\int_{\mathbb{R}_{+}}\frac{1}{(\sigma^{2})^{(k+\upsilon_{0})/2+1}}e^{-(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)/\sigma^{2}}\mathrm{d}(\sigma^{2})\right)e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}\mathrm{d}(\beta)$
$=\int_{\mathbb{R}^{q}}\frac{\Gamma((k+\upsilon_{0})/2)}{(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)^{(k+\upsilon_{0})/2}}e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}\mathrm{d}(\beta)$
$=\Gamma((k+\upsilon_{0})/2)\int_{\mathbb{R}^{q}}\frac{e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}}{(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)^{(k+\upsilon_{0})/2}}\mathrm{d}(\beta)$
$\leq\Gamma((k+\upsilon_{0})/2)\int_{B}\frac{e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}}{(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)^{(k+\upsilon_{0})/2}}\mathrm{d}(\beta)$
where $B\subset\mathbb{R}^{q}\times\mathbb{R}_{+}$ is a bounded set.
Thus we get that
$K$ $=\frac{1}{\int_{\mathbb{R}^{q}\times\mathbb{R}_{+}}g(\beta,\sigma^{2})\mathrm{d}(\beta,\sigma^{2})}\sup_{(\beta,\sigma^{2})\in\mathbb{R}\times\mathbb{R}_{+}}\frac{g(\beta,\sigma^{2})}{f_{\nu}(\beta,\sigma^{2})}$
$\leq\frac{1}{\Gamma((k+\upsilon_{0})/2)\int_{B}\frac{e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}}{(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)^{(k+\upsilon_{0})/2}}\mathrm{d}(\beta)}$
$\hskip 56.9055pt\frac{16\Gamma((k+\upsilon_{0})/2-2)(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{((Y-X\hat{\beta})^{T}(Y-X\hat{\beta}))^{2}\left(\upsilon_{0}c^{2}_{0}/2\right)^{(k+\upsilon_{0})/2-2}e^{2}}$
$=\frac{1}{\int_{B}\frac{e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}}{(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)^{(k+\upsilon_{0})/2}}\mathrm{d}(\beta)}$
$\hskip 56.9055pt\frac{16\Gamma((k+\upsilon_{0})/2-2)(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{\Gamma((k+\upsilon_{0})/2)((Y-X\hat{\beta})^{T}(Y-X\hat{\beta}))^{2}\left(\upsilon_{0}c^{2}_{0}/2\right)^{(k+\upsilon_{0})/2-2}e^{2}}$
So by Equation 3 combined with Theorem 3.2, for $r\geq p$
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq\left(\frac{C}{L}E[d(X_{n},Y_{n})^{r}]\right)^{1/r}$
Where $C=\frac{16\Gamma((k+\upsilon_{0})/2-2)(2\pi)^{q/2}\det(\Sigma_{\beta})^{1/2}}{\Gamma((k+\upsilon_{0})/2)((Y-X\hat{\beta})^{T}(Y-X\hat{\beta}))^{2}\left(\upsilon_{0}c^{2}_{0}/2\right)^{(k+\upsilon_{0})/2-2}e^{2}}$
and $L=\int_{B}\frac{e^{-\frac{1}{2}(\beta-\beta_{0})^{T}\Sigma_{\beta}^{-1}(\beta-\beta_{0})}}{(\upsilon_{0}c_{0}^{2}/2+(Y-X\beta)^{T}(Y-X\beta)/2)^{(k+\upsilon_{0})/2}}\mathrm{d}(\beta)$
Remark 3.
[…]